OpenStudy (anonymous):

What is the point-slope and standard forms of this? (3,-2)(6,2)

OpenStudy (karatechopper):

Point slope: (y-y1)=m(x-x1) Standard form: Ax+By=C

OpenStudy (karatechopper):

Btw, Welcome to OpenStudy :)

OpenStudy (anonymous):

thank you :)